# for functionals in the $\mathcal{A}$-free setting

### v1

@@ 5. The integrand obtained from minimum values on small cubes
Let $f\in\mathcal{F}_{\mathrm{qc}}$. Then, for a.e.  $x\in\mathbb{R}^{N}$ and every $\xi\in\mathbb{R}^{d}$, we have that
@@
from which (5.6) follows by taking the limit $k\to\infty$. Since $f\in\mathcal{F}_{\mathrm{qc}}$ we may assume, changing possibly $E$, that for every $x\in\mathbb{R}^{N}\setminus E$, the function $\xi\mapsto f(x,\xi)$ is $\mathcal{A}$-quasiconvex.
@@ 6. The $\Gamma$-limit obtained from minimum values on small cubes
Let $(f_{k})_{k\in\mathbb{N}}$ be a sequence in $\mathcal{F}$, let $f\in\mathcal{F}$, and let $F_{k}$ and $F$ be the corresponding functionals in $\mathcal{I}$ defined by (2.13). Assume that for every $D\in\mathcal{O}(\mathbb{R}^{N})$, the sequence $(F_{k}(\cdot,D))_{k\in\mathbb{N}}$ $\Gamma$-converges to $F(\cdot,D)$ with respect to the topology induced by $\|\cdot\|_{D}^{\mathcal{A}}$ on $L^{p}(D;\mathbb{R}^{d})$. Let $Q\subset\mathbb{R}^{N}$ be an open cube and let $\xi\in\mathbb{R}^{d}$. Then,
@@
Let $(f_{k})_{k\in\mathbb{N}}$ be a sequence in $\mathcal{F}$, let $(F_{k})_{k\in\mathbb{N}}$ be the corresponding sequence of functionals in $\mathcal{I}$ defined by (2.13), and let $(F_{k}^{\mathcal{A}})_{k\in\mathbb{N}}$ be the sequence obtained as in Definition 4.4. Suppose that
@@
Then, $f\in\mathcal{F}_{\mathrm{qc}}$ and the functionals $F$ and $F^{\mathcal{A}}$ introduced in (2.13) and Definition 4.4 satisfy the following properties:
@@
The equivalence between (b) and (c) is proved in Theorem 4.6. If these conditions are satisfied, then the function $f$ belongs to $\mathcal{F}_{\mathrm{qc}}$ by Corollary 3.4.
Next, we assume that (b) holds, and we show that $f$ satisfies (6.6). Fix $x\in\mathbb{R}^{N}$ and $\xi\in\mathbb{R}^{d}$. By Proposition 6.1, we have for every $\rho>0$ that
@@
By Theorem 5.3, we obtain (6.6) for a.e. $x\in\mathbb{R}^{N}$ and every $\xi\in\mathbb{R}^{d}$, concluding the proof of (a).
@@
By (6.6), this implies that $\hat{f}(x,\xi)=f(x,\xi)$ for a.e. $x\in\mathbb{R}^{N}$ and every $\xi\in\mathbb{R}^{d}$; hence, $f\in\mathcal{F}_{\mathrm{qc}}$ and $\widehat{F}=F$. Since the $\Gamma$-limit does not depend on the subsequence, we obtain (b) by the Urysohn property of $\Gamma$-convergence (see [11, Proposition 8.3]). ∎
@@
By the Hartman–Stampacchia Theorem [20, Lemma 3.1], there exists a unique function $v\in W^{1,q}_{m}(D;\mathbb{R}^{l})$ such that $T(v)=\mathcal{G}$; i.e.,
@@
It follows immediately from the definition of $M^{\eta}_{c}$ and from (2.9) that for every $\xi_{1}$ and $\xi_{2}\in\mathbb{R}^{d}$, we have
@@
Let $(f_{k})_{k\in\mathbb{N}}$ be a sequence in $\mathcal{F}$, let $f\in\mathcal{F}$, and let $F_{k}$ and $F$ be the corresponding functionals in $\mathcal{I}$ defined by (2.13). Assume that for every $D\in\mathcal{O}(\mathbb{R}^{N})$, the sequence $(F_{k}(\cdot,D))_{k\in\mathbb{N}}$ $\Gamma$-converges to $F(\cdot,D)$ with respect to the topology induced by $\|\cdot\|_{D}^{\mathcal{A}}$ on $L^{p}(D;\mathbb{R}^{d})$. Then, for every $\eta>0$, every cube $Q\subset\mathbb{R}^{N}$, and every $\xi\in\mathbb{R}^{d}$, we have
@@
Fix $\eta>0$. Since $\mathcal{A}_{Q}z_{k}\to 0$ in $W^{-1,p}(Q;\mathbb{R}^{l})$, we obtain from (6.28) that, for $k$ large enough, the functions $V_{k}$ belong to the set $\mathcal{V}^{\eta}(Q)$ introduced in Definition 6.5. Moreover, since $\operatorname{supp}z_{k}\subset\subset Q$ and $\int_{Q}z_{k}\,dx=0$, we have for $k$ large enough that the functions $z_{k}$ belong to the set $\mathcal{U}^{\eta}_{c}(Q)$ introduced in Definition 6.5. By (6.16), this implies that $M^{\eta}_{c}(f_{k},\xi,Q)\leq F_{k}(\xi+z_{k},Q)$ that, together with (6.27), yields
@@
$\frac{1}{c_{0}}\int_{Q}|\xi+u_{k}|^{p\,}dx-c_{0}|Q|\leq\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}F_{k}(\xi+u_{k},Q)<M^{\eta}_{c}(f_{k},\xi,Q)+\tfrac{1}{k}|Q|\leq\Big{(}c_{0}(1+|\xi|^{p})+1\Big{)}|Q|,$ (6.29)
where the first and last inequality follow from (2.10). These inequalities imply that $\|u_{k}\|_{L^{p}(Q;\mathbb{R}^{d})}^{p}<C|Q|$ for every $k\in\mathbb{N}$, which allows us to extract a subsequence of $(u_{k})_{k\in\mathbb{N}}$, not relabeled, that converges to some function $u$ weakly in $L^{p}(Q;\mathbb{R}^{d})$.
@@
Let $(f_{k})_{k\in\mathbb{N}}$ be a sequence in $\mathcal{F}$, let $(F_{k})_{k\in\mathbb{N}}$ be the corresponding sequence of functionals in $\mathcal{I}$ defined by (2.13), and let $(F_{k}^{\mathcal{A}})_{k\in\mathbb{N}}$ be the sequence obtained as in Definition 4.4. Suppose that
@@
Then, $f\in\mathcal{F}_{\mathrm{qc}}$ and the functionals $F$ and $F^{\mathcal{A}}$ introduced in (2.13) and Definition 4.4 satisfy the following properties:
@@
The equivalence between (b) and (c) is proved in Theorem 4.6. If these conditions are satisfied, then the function $f$ belongs to $\mathcal{F}_{\mathrm{qc}}$ by Corollary 3.4.
@@
$f(x,\xi)$ $=\lim_{\rho\to 0^{+}}\frac{M(f,\xi,Q_{\rho}(x))}{\rho^{N}}\leq\limsup_{\rho\to 0^{+}}\sup_{\eta>0}\liminf_{k\to\infty}\frac{M^{\eta}_{c}(f_{k},\xi,Q_{\rho}(x))}{\rho^{N}}$
$\leq\limsup_{\rho\to 0^{+}}\sup_{\eta>0}\limsup_{k\to\infty}\frac{M^{\eta}_{c}(f_{k},\xi,Q_{\rho}(x))}{\rho^{N}}\leq\limsup_{\rho\to 0^{+}}\frac{M_{c}(f,\xi,Q_{\rho}(x))}{\rho^{N}}=f(x,\xi).$
@@
By (6.31), this implies that $\hat{f}(x,\xi)=f(x,\xi)$ for a.e. $x\in\mathbb{R}^{N}$ and every $\xi\in\mathbb{R}^{d}$; hence, $f\in\mathcal{F}_{\mathrm{qc}}$ and $\widehat{F}=F$. Since the $\Gamma$-limit does not depend on the subsequence, we obtain (b) from the Urysohn property of $\Gamma$-convergence (see [11, Proposition 8.3]). ∎
@@ 7. Homogenization without periodicity assumptions
Throughout this section, we fix a function $f\in\mathcal{F}$. For every $\varepsilon>0$, we consider the functions $f_{\varepsilon}\in\mathcal{F}$ defined by
@@
for every $x\in\mathbb{R}^{N}$ and $\xi\in\mathbb{R}^{d}$, the functionals $F_{\varepsilon}\in\mathcal{I}$ associated with $f_{\varepsilon}$ by (2.13), and the corresponding $\mathcal{A}$-free functionals $F_{\varepsilon}^{\mathcal{A}}$ introduced in Definition 4.4.
The following theorem provides very general conditions on the function $f$ which ensure that there exists a function $f_{\mathrm{hom}}\in\mathcal{F}$, independent of $x$, such that for every $D\in\mathcal{O}(\mathbb{R}^{N})$, the family of functionals $(F_{\varepsilon}(\cdot,D))_{\varepsilon>0}$ $\Gamma$-converges as $\varepsilon\to 0^{+}$ to the functional $F_{\mathrm{hom}}(\cdot,D)$ corresponding to $f_{\mathrm{hom}}$. By this we mean that for every sequence $(\varepsilon_{k})_{k\in\mathbb{N}}$ of positive numbers converging to 0, the sequence $(F_{\varepsilon_{k}}(\cdot,D))_{k\in\mathbb{N}}$ $\Gamma$-converges to $F_{\mathrm{hom}}(\cdot,D)$.
@@
for every $\xi\in\mathbb{R}^{d}$. Then, $f_{\mathrm{hom}}\in\mathcal{F}_{\mathrm{qc}}$ and the following properties hold:
@@
Assume that $f\in\mathcal{F}$ can be written as
@@
where $f_{per},\,f_{comp}\in\mathcal{F}$ satisfy the two following properties:
$(a)\enspace$ $x\mapsto f_{per}(x,\xi)\hbox{ is }Q_{1}(0)\hbox{-periodic for every }\xi\in\mathbb{R}^{d};$ (7.4)
@@
where in the last equality we used the monotonicity of $M^{\eta}_{c}$ with respect to $\eta$. Then, by Theorem 6.9, the function $f_{\mathrm{hom}}$ belongs to $\mathcal{F}_{\mathrm{qc}}$ and (a) and (b) are satisfied. ∎

### v2

@@ 5. The integrand obtained from minimum values on small cubes
Let $f\in\mathcal{F}_{\mathrm{Lip}}\cap\mathcal{F}_{\mathrm{qc}}$. Then, for a.e.  $x\in\mathbb{R}^{N}$ and every $\xi\in\mathbb{R}^{d}$, we have that
@@
from which (5.6) follows by taking the limit $k\to\infty$. Since $f\in\mathcal{F}_{\mathrm{qc}}$ for every $x\in\mathbb{R}^{N}\setminus E$ the function $\xi\mapsto f(x,\xi)$ is $\mathcal{A}$-quasiconvex.
@@ 6. The $\Gamma$-limit obtained from minimum values on small cubes
Let $(f_{k})_{k\in\mathbb{N}}$ be a sequence in $\mathcal{F}_{\mathrm{Lip}}$, let $f\in\mathcal{F}$, and let $F_{k}$ and $F$ be the corresponding functionals in $\mathcal{I}$ defined by (2.14). Assume that for every $D\in\mathcal{O}(\mathbb{R}^{N})$, the sequence $(F_{k}(\cdot,D))_{k\in\mathbb{N}}$ $\Gamma$-converges to $F(\cdot,D)$ with respect to the topology induced by $\|\cdot\|_{D}^{\mathcal{A}}$ on $L^{p}(D;\mathbb{R}^{d})$. Let $Q\subset\mathbb{R}^{N}$ be an open cube and let $\xi\in\mathbb{R}^{d}$. Then,
@@
Let $(f_{k})_{k\in\mathbb{N}}$ be a sequence in $\mathcal{F}_{\mathrm{Lip}}$, let $(F_{k})_{k\in\mathbb{N}}$ be the corresponding sequence of functionals in $\mathcal{I}$ defined by (2.14), and let $(F_{k}^{\mathcal{A}})_{k\in\mathbb{N}}$ be the sequence obtained as in Definition 4.4. Suppose that
@@
Then, there exists $\hat{f}\in\mathcal{F}_{\mathrm{Lip}}\cap\mathcal{F}_{\mathrm{qc}}$ such that $f(x,\xi)=\hat{f}(x,\xi)$ for a.e. $x\in\mathbb{R}^{N}$ and every $\xi\in\mathbb{R}^{d}$, and the functionals $F$ and $F^{\mathcal{A}}$ introduced in (2.14) and Definition 4.4 satisfy the following properties:
@@
Since $F_{k}\in\mathcal{F}_{\mathrm{Lip}}$ for every $k\in\mathbb{N}$ by Remark 2.11, the equivalence between (b) and (c) is proved in Theorem 4.6. If these conditions are satisfied, then by Corollary 3.4 there exists $\hat{f}\in\mathcal{F}_{\mathrm{Lip}}\cap\mathcal{F}_{\mathrm{qc}}$ such that
$F(u,D):=\int_{D}\hat{f}(x,u(x))\,dx,\quad\text{for every }D\in\mathcal{O}(\mathbb{R}^{N})\text{ and }u\in L^{p}(D;\mathbb{R}^{d}),$
which implies $f(x,\xi)=\hat{f}(x,\xi)$ for a.e. $x\in\mathbb{R}^{N}$ and every $\xi\in\mathbb{R}^{d}$.
Next, we assume that (b) holds, and we show that $\hat{f}$ satisfies (6.6). Fix $x\in\mathbb{R}^{N}$ and $\xi\in\mathbb{R}^{d}$. By Proposition 6.1 for every $\rho>0$ we have that
@@
By Theorem 5.3, we obtain (6.6) with $\hat{f}$ for a.e. $x\in\mathbb{R}^{N}$ and every $\xi\in\mathbb{R}^{d}$, concluding the proof of (a).
@@
By (6.6), this implies that $\hat{f}(x,\xi)=f(x,\xi)$ for a.e. $x\in\mathbb{R}^{N}$ and every $\xi\in\mathbb{R}^{d}$; hence, $\widehat{F}=F$. Since the $\Gamma$-limit does not depend on the subsequence, we obtain (b) by the Urysohn property of $\Gamma$-convergence (see [12, Proposition 8.3]). ∎
@@
By the Hartman–Stampacchia Theorem [22, Lemma 3.1], there exists a unique function $v\in W^{1,q}_{m}(D;\mathbb{R}^{l})$ such that $T(v)=\mathcal{G}$; i.e.,
@@
If $f\in\mathcal{F}_{\mathrm{Lip}}$, it follows immediately from the definition of $M^{\eta}_{c}$ and from (2.12) that for every $\xi_{1}$ and $\xi_{2}\in\mathbb{R}^{d}$, we have
@@
Let $(f_{k})_{k\in\mathbb{N}}$ be a sequence in $\mathcal{F}_{\mathrm{Lip}}$, let $f\in\mathcal{F}$, and let $F_{k}$ and $F$ be the corresponding functionals in $\mathcal{I}$ defined by (2.14). Assume that for every $D\in\mathcal{O}(\mathbb{R}^{N})$, the sequence $(F_{k}(\cdot,D))_{k\in\mathbb{N}}$ $\Gamma$-converges to $F(\cdot,D)$ with respect to the topology induced by $\|\cdot\|_{D}^{\mathcal{A}}$ on $L^{p}(D;\mathbb{R}^{d})$. Then, for every $\eta>0$, every cube $Q\subset\mathbb{R}^{N}$, and every $\xi\in\mathbb{R}^{d}$, we have
@@
Fix $\eta>0$. Since $\mathcal{A}_{Q}z_{k}\to 0$ in $W^{-1,p}(Q;\mathbb{R}^{l})$, we obtain from (6.28) that, for $k$ large enough, the functions $V_{k}$ belong to the set $\mathcal{V}^{\eta}(Q)$ introduced in Definition 6.5. Moreover, since $\operatorname{supp}z_{k}\subset\subset Q$ and $\int_{Q}z_{k}\,dx=0$, we have for $k$ large enough that the functions $z_{k}$ belong to the set $\mathcal{U}^{\eta}_{c}(Q)$ introduced in Definition 6.5. By (6.16), this implies that $M^{\eta}_{c}(f_{k},\xi,Q)\leq F_{k}(\xi+z_{k},Q)$. Together with (6.27), this yields
@@
$\frac{1}{c_{0}}\int_{Q}|\xi+u_{k}|^{p\,}dx-c_{0}|Q|\leq F_{k}(\xi+u_{k},Q)<M^{\eta}_{c}(f_{k},\xi,Q)+\tfrac{1}{k}|Q|\leq\Big(c_{0}(1+|\xi|^{p})+1\Big)|Q|,$ (6.29)
where the first and last inequality follow from (a) in Definition 2.9. These inequalities imply that $\|u_{k}\|_{L^{p}(Q;\mathbb{R}^{d})}^{p}<C|Q|$ for every $k\in\mathbb{N}$, which allows us to extract a subsequence of $(u_{k})_{k\in\mathbb{N}}$, not relabeled, that converges to some function $u$ weakly in $L^{p}(Q;\mathbb{R}^{d})$.
@@
Let $(f_{k})_{k\in\mathbb{N}}$ be a sequence in $\mathcal{F}_{\mathrm{Lip}}$, let $(F_{k})_{k\in\mathbb{N}}$ be the corresponding sequence of functionals in $\mathcal{I}$ defined by (2.14), and let $(F_{k}^{\mathcal{A}})_{k\in\mathbb{N}}$ be the sequence obtained as in Definition 4.4. Suppose that
@@
Then, there exists $\hat{f}\in\mathcal{F}_{\mathrm{Lip}}\cap\mathcal{F}_{\mathrm{qc}}$ such that $f(x,\xi)=\hat{f}(x,\xi)$ for a.e. $x\in\mathbb{R}^{N}$ and every $\xi\in\mathbb{R}^{d}$, and the functionals $F$ and $F^{\mathcal{A}}$ introduced in (2.14) and Definition 4.4 satisfy the following properties:
@@
Since $F_{k}\in\mathcal{F}_{\mathrm{Lip}}$ for every $k\in\mathbb{N}$ by Remark 2.11, the equivalence between (b) and (c) is proved in Theorem 4.6. If these conditions are satisfied, then by Corollary 3.4 there exists $\hat{f}\in\mathcal{F}_{\mathrm{Lip}}\cap\mathcal{F}_{\mathrm{qc}}$ such that
$F(u,D):=\int_{D}\hat{f}(x,u(x))\,dx,\quad\text{for every }D\in\mathcal{O}(\mathbb{R}^{N})\text{ and }u\in L^{p}(D;\mathbb{R}^{d}),$
which implies that $f(x,\xi)=\hat{f}(x,\xi)$ for a.e. $x\in\mathbb{R}^{N}$ and every $\xi\in\mathbb{R}^{d}$.
@@
$\hat{f}(x,\xi)=\lim_{\rho\to 0^{+}}\frac{M(\hat{f},\xi,Q_{\rho}(x))}{\rho^{N}}=\lim_{\rho\to 0^{+}}\frac{M(f,\xi,Q_{\rho}(x))}{\rho^{N}}\leq\limsup_{\rho\to 0^{+}}\sup_{\eta>0}\liminf_{k\to\infty}\frac{M^{\eta}_{c}(f_{k},\xi,Q_{\rho}(x))}{\rho^{N}}$
$\leq\limsup_{\rho\to 0^{+}}\sup_{\eta>0}\limsup_{k\to\infty}\frac{M^{\eta}_{c}(f_{k},\xi,Q_{\rho}(x))}{\rho^{N}}\leq\limsup_{\rho\to 0^{+}}\frac{M_{c}(f,\xi,Q_{\rho}(x))}{\rho^{N}}=\limsup_{\rho\to 0^{+}}\frac{M_{c}(\hat{f},\xi,Q_{\rho}(x))}{\rho^{N}}=\hat{f}(x,\xi).$
@@
By (6.31), this implies that $\hat{f}(x,\xi)=f(x,\xi)$ for a.e. $x\in\mathbb{R}^{N}$ and every $\xi\in\mathbb{R}^{d}$; hence, $f\in\mathcal{F}_{\mathrm{qc}}$ and $\widehat{F}=F$. Since the $\Gamma$-limit does not depend on the subsequence, we obtain (b) from the Urysohn property of $\Gamma$-convergence (see [12, Proposition 8.3]). ∎
@@ 7. Homogenization without periodicity assumptions
Throughout this section, we fix a function $f\in\mathcal{F}_{\mathrm{Lip}}$. We observe that if the vector space $\mathrm{span}(\Lambda)$ generated by the wave cone $\Lambda$ coincides with $\mathbb{R}^{d}$ then every $f\in\mathcal{F}_{\mathrm{qc}}$ belongs to $\mathcal{F}_{\mathrm{Lip}}$ by Remark 2.8. For every $\varepsilon>0$, we consider the functions $f_{\varepsilon}\in\mathcal{F}$ defined by
@@
for every $x\in\mathbb{R}^{N}$ and $\xi\in\mathbb{R}^{d}$, the functionals $F_{\varepsilon}\in\mathcal{I}$ associated with $f_{\varepsilon}$ by (2.14), and the corresponding $\mathcal{A}$-free functionals $F_{\varepsilon}^{\mathcal{A}}$ introduced in Definition 4.4. Note that $f_{\varepsilon}\in\mathcal{F}_{\mathrm{Lip}}$ for every $\varepsilon>0$.
The following theorem provides very general conditions on the function $f$ which ensure that there exists a function $f_{\mathrm{hom}}\in\mathcal{F}$, independent of $x$, such that for every $D\in\mathcal{O}(\mathbb{R}^{N})$, the family of functionals $(F_{\varepsilon}(\cdot,D))_{\varepsilon>0}$ $\Gamma$-converges as $\varepsilon\to 0^{+}$ to the functional $F_{\mathrm{hom}}(\cdot,D)$ corresponding to $f_{\mathrm{hom}}$. By this we mean that, for every sequence $(\varepsilon_{k})_{k\in\mathbb{N}}$ of positive numbers converging to 0, the sequence $(F_{\varepsilon_{k}}(\cdot,D))_{k\in\mathbb{N}}$ $\Gamma$-converges to $F_{\mathrm{hom}}(\cdot,D)$.
@@
for every $\xi\in\mathbb{R}^{d}$. Then, $f_{\mathrm{hom}}\in\mathcal{F}_{\mathrm{Lip}}\cap\mathcal{F}_{\mathrm{qc}}$ and the following properties hold:
@@
Assume that $f$ can be written as
@@
where $f_{per},\,f_{comp}\in\mathcal{F}_{\mathrm{Lip}}$ satisfy the two following properties:
$(a)\kern 5.0pt$ $x\mapsto f_{per}(x,\xi)\hbox{ is }Q_{1}(0)\hbox{-periodic for every }\xi\in\mathbb{R}^{d};$ (7.4)
@@
where in the last equality we used the monotonicity of $M^{\eta}_{c}$ with respect to $\eta$. Then, by Theorem 6.9, the function $f_{\mathrm{hom}}$ belongs to $\mathcal{F}_{\mathrm{Lip}}\cap\mathcal{F}_{\mathrm{qc}}$ and (a) and (b) are satisfied. ∎